# Unifying F1TENTH Autonomous Racing: Survey, Methods and Benchmarks

### v1

@@ Abstract
The F1TENTH autonomous racing platform, consisting of 1:10 scale RC cars, has evolved into a leading research platform. The many publications and real-world competitions span many domains, from classical path planning to novel learning-based algorithms. Consequently, the field is wide and disjointed, hindering direct comparison of methods and making it difficult to assess the state-of-the-art. Therefore, we aim to unify the field by surveying current approaches, describing common methods and providing benchmark results to facilitate clear comparison and establish a baseline for future work. We survey current work in F1TENTH racing in the classical and learning categories, explaining the different solution approaches. We describe particle filter localisation, trajectory optimisation and tracking, model predictive contouring control (MPCC), follow-the-gap and end-to-end reinforcement learning. We provide an open-source evaluation of benchmark methods and investigate overlooked factors of control frequency and localisation accuracy for classical methods and reward signal and training map for learning methods. The evaluation shows that the optimisation and tracking method achieves the fastest lap times, followed by the MPCC planner. Finally, our work identifies and outlines the relevant research aspects to help motivate future work in the F1TENTH domain.
@@ I Introduction
The growing field of F1TENTH autonomous racing offers an exciting testbed for cutting-edge robotics and autonomous vehicle research [1]. The domain encompasses diverse aspects of autonomous systems, including perception, planning, control and learning [2]. However, the breath of F1TENTH research has left the field divided and fragmented, with researchers often focusing on specific subproblems in isolation. A particular divide exists between classical and machine learning approaches. This fragmentation hinders comparison between methods, making it difficult to assess algorithmic advancement. This paper aims to unify F1TENTH research by providing a comprehensive overview of the field and introducing a consistent perspective to unite different disciplines.
Autonomous racing presents a uniquely demanding testbed for the advancement of autonomous robotics methods. The racing problem is to select control actions (speed and steering) based on raw sensor data (LiDAR) that move the vehicle around the track as quickly as possible. The nature of racing provides a difficult challenge due to non-linear tyre dynamics, unstructured sensor data and real-time computing requirements [3]. A key challenge is the fundamental conflict between maximising performance and ensuring safety [4]. Higher speeds increase operational risk while prioritizing safety sacrifices competitive advantage. These difficulties promote the improvement of high-performance, safe, and efficient autonomous vehicle algorithms.
The F1TENTH platform is ideal for research because it enables rapid prototyping of algorithms, provides a competitive environment and simplifies software-to-hardware deployment [5]. F1TENTH vehicles, shown in Fig. 1, are equipped with a 2D-LiDAR for sensing, an onboard computer for processing, and actuator controllers for the motors [1]. The smaller size of F1TENTH cars makes them cost-effective and safe for developing new high-performance algorithms where crashes are inevitable. Finally, it provides a fun topic for students worldwide [6], and competitions encourage novel research.
@@
We unify the field of F1TENTH racing by creating a common solution taxonomy through the following contributions:
  • Survey advances in classical and learning-based approaches for F1TENTH racing in Section II.
  • Describe common classical and reinforcement learning baseline methods in Section III.
  • Provide benchmark results (with open-source software [^1] Code available at: https://github.com/BDEvan5/f1tenth_benchmarks) to accelerate future work in Section IV.
  • Motivate promising research directions for future work on F1TENTH cars in Section V.
@@ II-A Classical Approaches
Classical approaches to autonomous racing use estimation, optimisation and control systems to calculate control commands. Typically, estimation is used to localise the vehicle on the map, optimisation is used for planning, and tracking algorithms are used to follow the trajectory. Fig. 2 shows how classical pipelines use localisation with either offline planning and online control or online planning and control. We study works related to each of these functions.
Perception: Perception is the task of building a map and localising the vehicle on the map. While no methods of map building, i.e., simultaneous localization and mapping (SLAM), have been built explicitly for F1TENTH racing, the ROS SLAM toolbox (using graph SLAM) is commonly used [7, 8]. Walsh et al. [9] implement a computationally efficient method for particle filter localization. Their approach, and open-source repo, is the standard localization method used in many other approaches [10, 11]. An extension of their approach has improved the original implementation by using a higher-order motion model [12]. A different approach using invertible neural networks to learn a localisation policy boasted improved computational performance [13]. However, the particle filter remains the most commonly used method due to its simplicity and robustness.
Offline trajectory planning: Offline trajectory planning uses an optimisation algorithm and map of the track to generate an optimal set of waypoints. The baseline planning method for most works uses a minimum curvature or minimum time trajectory planning formulation from Heilmeier et al. [14], which formulates the trajectory planning problem as a quadratic programming problem. Other approaches have turned to genetic programming algorithms to jointly optimise the trajectory and the controller [10, 15], but these can lead to simulator exploitation and unrealistic behaviour [16]. Heilmeier et al.’s method is the most popular because it is efficient to compute, robust to different tracks, open-source accessible and provides high-quality results [17, 18].
Online trajectory planning: At each planning step, online trajectory planning approaches generate a set of control inputs and a corresponding trajectory for a receding horizon [19]. Model predictive contouring control (MPCC) has been used to optimise a trajectory that maximises progress along the centre line [20]. Since online planning algorithms replan at each timestep, they can more easily integrate obstacle avoidance and head-to-head racing strategies [20, 21]. While online planning algorithms, such as MPCC, can plan accurate, high-performance paths, they are limited by requiring accurate vehicle models and system identification [22]. Additionally, due to the complexity of implementing optimisation routines, there is little comparison between MPC formulations (such as between different dynamics models) or other methods.
@@
Control: The controller generates actuator commands for speed and steering that cause the vehicle to track the reference trajectory. The most common control algorithm is pure pursuit, a geometric path tracking algorithm based on the vehicle dynamics [23]. The method has been advanced by using a speed-dependant lookahead distance [11], and integrating the tyre dynamics into the algorithm to improve tracking at high speeds [18]. Model predictive control (MPC) approaches have also been used for control [24, 25], with the advantage that they can perform online system identification [26], thus updating the model as more data is collected [27]. Despite these advantages, the pure pursuit algorithm remains the most common control algorithm due to its robustness and simplicity [10, 5, 28].
@@
Classical O’Kelly et at. (2020) [10, 16] Genetic optimisation Offline planing F1TENTH Gym
Gradient-based trajectory optimisation [14]
@@
Curinga (2018) [19] MPCC Online planning Custom
Optimisation & tracking
@@
Jain et al. (2020) [27] MPC with gaussian process Control F1TENTH Gym
MPC
@@
Wang et al. (2021) [26] Data-driven MPC Control Gazebo
MPC, adaptive pure pursuit
Nagy et al. (2023) [24] Adaptive MPC Control F1TENTH Gym
Pure pursuit
Sukhil and Behl (2021) [11] Adaptive pure pursuit Control F1Tenth.dev
Pure pursuit
Becker et al. (2023) [18] Model- & acceleration-based pure pursuit Control
Pure pursuit
Sezer et at. (2012) [29] Follow-the-gap Mapless Custom
APF, A*
@@
Learning Hamilton et al. (2022) [31] DRL vs IL End-to-end Gazebo
IL, SAC, DDPG
Brunnbauer et al. (2022) [32] Model-based DRL End-to-end PyBullet
Model-free DRL
Bosello et al. (2022) [33] Generalisable DRL End-to-end Rviz
DRL, Dreamer, FTG
Sun et al. (2023)[34] IL algorithm comparison End-to-end F1TENTH Gym
IL, PPO, DAgger, EIL
Sun et al. (2023) [35] New IL algorithm End-to-end F1TENTH Gym
IL, experts
Evans et al. (2023) [36] Trajectory-aided learning reward End-to-end F1TENTH Gym
Cross-track & heading
Evans et al. (2023) [7] Comparing DRL architectures Comparison F1TENTH Gym
Optimisation & tracking
Tătulea-Codrean et al. (2020) [37] IL to replace MPC Learned planning Gazebo
MPC
Ghignone et al. (2022) [22] Trajectory-conditioned learning Learned planning F1TENTH Gym
MPC
Dwivedi et al. (2022) [28] Plan-assisted DRL Learned planning PyBullet
Dreamer, MPO
Trumpp et al. (2023) [38] RPL for high-speed racing RPL F1TENTH Gym
Optimisation & tracking
Evans et al. (2021) [39, 40] Static obstacle avoidance RPL Custom
FTG, end-to-end DRL
Zhang et al. (2022) [41] APF-based RPL RPL PyBullet
FTG, MPC, DRL, Dreamer
Trumpp et al. (2024) [42] Mapless RPL racing RPL F1TENTH Gym
APF, disparity extender
Ivanov et al. (2020) [43] Formal neural network verification Safe learning Custom
DRL
Musau et al. (2022) [44] Reachability Theory safety Safe learning Gazebo
IL, DRL
Evans et al. (2023) [45, 46] Viability Theory safety Safe learning F1TENTH Gym
DRL, pure pursuit
Mapless Methods: A commonly used mapless approach for F1TENTH racing is the Follow-the-Gap (FTG) controller. Proposed by Sezer et at. [29], this is a reactive method based on identifying gaps in the car’s LiDAR signal. After finding the gaps, the car’s steering angle is directly defined as the angle in direction or the widest or furthest. An extension to FTG is the disparity extender [30]. In this approach, before detecting the gaps in the LiDAR signal, all disparities in the LiDAR signal are identified. The distance readings of the LiDAR signal adjacent in some distance to the disparity are set to the minimum distance in each interval. Typically, the interval is chosen as the vehicle width plus a safety margin, thus avoiding trajectories that cut obstacles closely.
@@ II-B Learning Approaches
End-to-end learning uses a neural network to replace the entire racing pipeline. In F1TENTH racing, common inputs for end-to-end agents are single or multiple downsampled 2D LiDAR scans and the vehicle speed [34, 32]. Many studies have addressed constant speed racing (driving) [31, 46], and implemented their approaches on physical vehicles at low, constant speeds [33, 32]. End-to-end approaches generalise well to unseen tracks [33] and transfer to physical vehicles [7]. A major challenge for end-to-end learning agents is reliable performance at high speed [36].
@@ III Racing Methods
We describe the most common racing algorithms from each category to provide an overview of different solution approaches. We describe the particle filter algorithm used for localisation, a trajectory planning and control strategy, the model predictive contouring algorithm, the follow-the-gap method and end-to-end deep reinforcement learning.
@@ III-A Particle Filter Algorithm
The particle filter (PF) algorithm uses the control input and the 2D LiDAR scan to estimate the vehicle’s pose on the map [9]. The algorithm is a recursive Bayesian state estimator designed for non-parametric distributions by representing the belief of the vehicle’s pose by a set of particles [50].
@@ IV Benchmark Evaluation
Methodology: We use the open-source F1TENTH autonomous vehicle simulator [5]. The vehicle is represented by the single-track bicycle model (described in Appendix A), which describes the state in terms of position, speed, steering angle, yaw, yaw rate, and slip angle [53]. The 2D LiDAR scanner is simulated with a ray-casting algorithm that measures the distance to the track boundary for each beam. The simulator and all the tests are written in Python and run on an Ubuntu computer. The simulator dynamics are updated at a rate of 100 Hz in all tests, and unless otherwise stated, the planning frequency is 25 Hz.
@@ IV-C Benchmark Results
AUT
ESP
GBR
MCO
Opti. & tracking
16.79 35.92 31.24 28.08
MPCC
16.87
39.13
35.40
31.53
Follow-the-gap
19.10
45.78
39.34
34.99
End-to-end
19.94
46.37
40.22
34.93
@@ V-A Future Work
Vision-based racing: Almost all racing methods use LiDAR as the main input modality. While LiDAR’s are convenient since they provide geometric information about the surrounding environment, they are expensive and poorly suited to real-world applications. In contrast, cameras are cheaper and contain more expressive information. Therefore, using cameras for high-performance control should be investigated. Caie et al. [54] started work in this direction with end-to-end neural networks; however, their solution is limited to slow speeds and neglects the key racing challenges.
@@ Appendix A Single-Track Model
The inputs into the system $u_{1}$ and $u_{2}$ are the steering angle velocity $v_{\delta}$, and longitudinal acceleration $a$. Using the parameters of vehicle mass $m$, cornering stiffness $C_{s}$ and wheelbase length $l$ for the front ${}_{f}$ and rear ${}_{r}$, the height of the centre of gravity $h_{\textbf{cg}}$, and gravitational constant $g$, the state derivatives are given by,

### v2

@@ Abstract
The F1TENTH autonomous driving platform, consisting of 1:10-scale remote-controlled cars, has evolved into a well-established education and research platform. The many publications and real-world competitions span many domains, from classical path planning to novel learning-based algorithms. Consequently, the field is wide and disjointed, hindering direct comparison of developed methods and making it difficult to assess the state-of-the-art. Therefore, we aim to unify the field by surveying current approaches, describing common methods, and providing benchmark results to facilitate clear comparisons and establish a baseline for future work. This research aims to survey past and current work with F1TENTH vehicles in the classical and learning categories and explain the different solution approaches. We describe particle filter localisation, trajectory optimisation and tracking, model predictive contouring control, follow-the-gap, and end-to-end reinforcement learning. We provide an open-source evaluation of benchmark methods and investigate overlooked factors of control frequency and localisation accuracy for classical methods as well as reward signal and training map for learning methods. The evaluation shows that the optimisation and tracking method achieves the fastest lap times, followed by the online planning approach. Finally, our work identifies and outlines the relevant research aspects to help motivate future work in the F1TENTH domain.
@@ I Introduction
The growing field of F1TENTH autonomous driving offers an exciting testbed for cutting-edge robotics and autonomous vehicle research [1]. The domain encompasses diverse aspects of autonomous systems, including perception, planning, control, and learning, and especially focuses on adversarial and high-speed environments [2]. Autonomous racing presents a uniquely demanding testbed for advancing autonomous robotics methods. The racing problem is to select control actions (speed and steering) based on raw sensor data (LiDAR) that move the vehicle around the track as quickly as possible. The nature of racing provides a difficult challenge due to non-linear tyre dynamics, unstructured sensor data and real-time computing requirements [3]. A key challenge is the fundamental conflict between maximising performance and ensuring safety [4]. Higher speeds increase operational risk while prioritizing safety, which sacrifices competitive advantage. These difficulties promote the improvement of high-performance, safe, and efficient autonomous vehicle algorithms.
The F1TENTH platform is ideal for research because it enables rapid prototyping of autonomy algorithms, provides a competitive environment and simplifies software-to-hardware deployment [5]. F1TENTH vehicles, shown in Fig. 1, are equipped with a 2D-LiDAR for sensing, an onboard computer for processing, and motor controllers [1]. The smaller size of F1TENTH cars makes them cost-effective and safe for developing new high-performance algorithms where crashes are inevitable. Finally, it provides a fun topic for students worldwide [6], and competitions encourage novel research.
However, the breath of F1TENTH research has left the field divided and fragmented, with researchers often focusing on specific subproblems in isolation. A particular divide exists between classical and machine learning approaches. This fragmentation hinders comparison between methods, making it difficult to assess algorithmic advancement. This paper aims to unify F1TENTH research by providing a comprehensive overview of the field and introducing a consistent perspective to unite different disciplines.
@@
We achieve this by creating a common solution taxonomy through the following contributions:
  • We provide a literature survey of advances in classical and learning-based approaches that leverage the F1TENTH vehicle.
  • We describe common classical and reinforcement learning baseline methods
  • We provide benchmark results with open-source software [^1] Code available at: https://github.com/BDEvan5/f1tenth_benchmarks to accelerate future work.
  • We describe promising research directions for future work on F1TENTH cars to improve current methods and increase the applicability of approaches.
@@ II-A Classical Approaches
Classical approaches to autonomous racing use a sense-plan-act structure, where the problem is split into distinct sub-problems. Typically, estimation is used to localise the vehicle on the map, optimisation is used for planning, and tracking algorithms are used to follow the trajectory. Fig. 2 shows how classical pipelines use localisation with either offline planning and online control or online planning and control. We study works related to each of these functions.
Perception: Perception is the task of building a map, localising the vehicle on the map, and perceiving objects. While no methods of map building, i.e., simultaneous localisation and mapping (SLAM), have been built explicitly for F1TENTH racing, the ROS SLAM toolbox (using graph SLAM) is commonly used [7, 8]. Walsh et al. [9] implement a computationally efficient method for particle filter localisation. Their approach, and open-source repo, is the standard localisation method used in many other approaches [10, 11]. An extension of their approach has improved the original implementation by using a higher-order motion model [12]. A different approach using invertible neural networks to learn a localisation policy boasted improved computational performance [13]. However, the particle filter remains the most commonly used method due to its simplicity and robustness.
Offline trajectory planning (global planning): Offline trajectory planning uses an optimisation algorithm and map of the track to generate an optimal set of waypoints. The baseline planning method for most works uses a shortest path, minimum curvature or minimum time trajectory planning formulation from Heilmeier et al. [14], which formulates the trajectory planning problem as a quadratic programming problem. Other approaches have turned to genetic programming algorithms to jointly optimise the trajectory and the controller [10, 15], but these can lead to simulator exploitation and unrealistic behaviour [16]. Heilmeier et al.’s method is the most popular because it is efficient to compute, robust to different tracks, open-source accessible and provides high-quality results [17, 18].
Online trajectory planning (local planning): At each planning step, online trajectory planning approaches generate a set of control inputs and a corresponding trajectory for a receding horizon [19]. Model predictive contouring control (MPCC) has been used to optimise a trajectory that maximises progress along the centre line [20]. Since online planning algorithms replan at each timestep, they can more easily integrate obstacle avoidance and head-to-head racing strategies [20, 21]. While online planning algorithms, such as MPCC, can plan accurate, high-performance paths, they are limited by requiring accurate vehicle models and system identification [22]. Additionally, due to the complexity of implementing optimisation routines, there is little comparison between Model Predictive Control (MPC) formulations (such as between different dynamics models) or other methods.
@@
Control: The controller generates actuator commands for speed and steering that cause the vehicle to track the reference trajectory. The most common control algorithm is pure pursuit, a geometric path tracking algorithm based on the vehicle dynamics [23]. The method has been advanced by using a speed-dependant lookahead distance [11], and integrating the tyre dynamics into the algorithm to improve tracking at high speeds [18]. While MPC is often used as a local planning method, it can also be used as a control approach [24, 25]. When using MPC for control, the objective is to track the global plan as accurately as possible. Using MPC for control has the advantage that the system identification can be performed online [26], which means that as the model accuracy can be continuously improved as more data is collected [27]. Despite these advantages, the pure pursuit algorithm remains the most common control algorithm due to its robustness and simplicity [10, 5, 28].
@@
Classical O’Kelly et at. (2020) [10, 16] Genetic optimisation Offline planing F1TENTH Gym Gradient-based trajectory optimisation [14]
@@
Curinga (2018) [19] MPCC Online planning Custom Optimisation & tracking
@@
Jain et al. (2020) [27] MPC with gaussian process Control F1TENTH Gym MPC
@@
Wang et al. (2021) [26] Data-driven MPC Control Gazebo MPC, adaptive pure pursuit
Nagy et al. (2023) [24] Adaptive MPC Control F1TENTH Gym Pure pursuit
Sukhil and Behl (2021) [11] Adaptive pure pursuit Control F1Tenth.dev Pure pursuit
Becker et al. (2023) [18] Model- & acceleration-based pure pursuit Control Pure pursuit
Sezer et at. (2012) [29] Follow-the-gap Mapless Custom APF, A*
@@
Learning Hamilton et al. (2022) [31] DRL vs IL End-to-end Gazebo IL, SAC, DDPG
Brunnbauer et al. (2022) [32] Model-based DRL End-to-end PyBullet Model-free DRL
Bosello et al. (2022) [33] Generalisable DRL End-to-end Rviz DRL, Dreamer, FTG
Sun et al. (2023)[34] IL algorithm comparison End-to-end F1TENTH Gym IL, PPO, DAgger, EIL
Sun et al. (2023) [35] New IL algorithm End-to-end F1TENTH Gym IL, experts
Evans et al. (2023) [36] Trajectory-aided learning reward End-to-end F1TENTH Gym Cross-track & heading
Evans et al. (2023) [7] Comparing DRL architectures Comparison F1TENTH Gym Optimisation & tracking
Tătulea-Codrean et al. (2020) [37] IL to replace MPC Learned planning Gazebo MPC
Ghignone et al. (2022) [22] Trajectory-conditioned learning Learned planning F1TENTH Gym MPC
Dwivedi et al. (2022) [28] Plan-assisted DRL Learned planning PyBullet Dreamer, MPO
Trumpp et al. (2023) [38] RPL for high-speed racing RPL F1TENTH Gym Optimisation & tracking
Evans et al. (2021) [39, 40] Static obstacle avoidance RPL Custom FTG, end-to-end DRL
Zhang et al. (2022) [41] APF-based RPL RPL PyBullet FTG, MPC, DRL, Dreamer
Trumpp et al. (2024) [42] Mapless RPL racing RPL F1TENTH Gym APF, disparity extender
Ivanov et al. (2020) [43] Formal neural network verification Safe learning Custom DRL
Musau et al. (2022) [44] Reachability Theory safety Safe learning Gazebo IL, DRL
Evans et al. (2023) [45, 46] Viability Theory safety Safe learning F1TENTH Gym DRL, pure pursuit
Mapless Methods: A commonly used mapless approach for F1TENTH racing is the Follow-the-Gap (FTG) controller. Proposed by Sezer et at. [29], this is a reactive method based on identifying gaps in the car’s LiDAR signal. After identifying all gaps in the current LiDAR scan, the car’s steering angle is directly defined as the angle in direction of the widest or furthest gap. An extension to FTG is the disparity extender [30]. In this approach, before detecting the gaps in the LiDAR signal, all disparities in the LiDAR signal are identified. The distance readings of the LiDAR signal adjacent in some distance to the disparity are set to the minimum distance in each interval. Typically, the interval is chosen as the vehicle width plus a safety margin, thus avoiding trajectories that cut obstacles closely.
@@ II-B Learning Approaches
End-to-end learning uses a neural network to replace the entire racing pipeline. In F1TENTH racing, common inputs for end-to-end agents are single or multiple downsampled 2D-LiDAR scans and the vehicle speed [34, 32]. Many studies have addressed constant speed racing (driving) [31, 46], and implemented their approaches on physical vehicles at low, constant speeds [33, 32]. End-to-end approaches generalise well to unseen tracks [33] and transfer to physical vehicles [7]. A major challenge for end-to-end learning agents is reliable performance at high speed [36].
@@ III Racing Methods
Currently, many methods for driving and racing autonomously with the F1TENTH vehicle are available. To unify this, we describe here the most common racing algorithms from each category to provide an overview of different solution approaches. We illustrate the particle filter algorithm used for localisation, a trajectory planning and control strategy, the model predictive contouring algorithm, the follow-the-gap method and end-to-end deep reinforcement learning.
@@ III-A Particle Filter Algorithm
The particle filter (PF) algorithm uses the control input and the 2D-LiDAR scan to estimate the vehicle’s pose on the map [9]. The algorithm is a recursive Bayesian state estimator designed for non-parametric distributions by representing the belief of the vehicle’s pose by a set of particles [50].
@@ IV Benchmark Evaluation
Methodology: We use the open-source F1TENTH autonomous vehicle simulator [5]. The vehicle is represented by the single-track bicycle model (described in Appendix A), which describes the state in terms of position, speed, steering angle, yaw, yaw rate, and slip angle [53]. The 2D-LiDAR scanner is simulated with a ray-casting algorithm that measures the distance to the track boundary for each beam. The simulator and all the tests are written in Python and run on an Ubuntu computer. The simulator dynamics are updated at a rate of 100 Hz in all tests, and unless otherwise stated, the planning frequency is 25 Hz.
@@ IV-C Benchmark Results
AUT ESP GBR MCO
Opti. & tracking 16.79 35.92 31.24 28.08
MPCC 16.87 39.13 35.40 31.53
Follow-the-gap 19.10 45.78 39.34 34.99
End-to-end 19.94 46.37 40.22 34.93
@@ V-A Future Work
Vision-based racing: Almost all racing methods use LiDAR as the main input modality. While LiDAR’s are convenient since they provide geometric information about the surrounding environment, they are expensive and poorly suited to real-world applications. In contrast, cameras are cheaper and contain more expressive information. Therefore, using cameras for high-performance control should be investigated. Cai et al. [54] started work in this direction with end-to-end neural networks; however, their solution is limited to slow speeds and neglects the key racing challenges.
3D-Aware Racing: While the works in this paper focus on racing in 2-dimensions the world exists in 3-dimensions. An interesting initial direction is 3D simulators that can accurately simulate 3D-LiDAR and visual camera input. Additional research areas are 3D mapping, object detection and localisation. Solutions focusing on 3-dimensions will improve the applicability of many methods to real-world robotics and autonomous driving pipelines.
@@ Appendix A Single-Track Model
The inputs into the system $u_{1}$ and $u_{2}$ are the steering angle velocity $v_{\delta}$, and longitudinal acceleration $a$. Using the parameters of vehicle mass $m$, cornering stiffness $C_{s}$ and wheelbase length $l$ for the front f and rear r, the height of the centre of gravity $h_{\textbf{cg}}$, and gravitational constant $g$, the state derivatives are given by,